# RLeXplore: Accelerating Research in Intrinsically-Motivated Reinforcement Learning

### v1

@@ 6 Conclusion
As the RL community continues to tackle increasingly complex tasks, the adoption of intrinsic rewards promises to play a pivotal role in improving agent autonomy and adaptability in diverse environments. In conclusion, RL agents can benefit significantly from intrinsic rewards. These enable the discovery of emergent behaviours that can make RL agents more autonomous and proficient in solving complex sparse-reward tasks. Our work introduces RLeXplore, a comprehensive open-source repository that not only implements state-of-the-art intrinsic rewards but also provides a systematic evaluation framework for understanding their impact on agent performance. We design RLeXplore to accelerate research in intrinsically-motivated RL and support fair comparsisons with well-stablished baseliens.
@@
Our results show that with RLeXplore, RL agents can learn emergent behaviours autonomously, solving multiple levels of SuperMarioBros without task rewards. Additionally, we show that intrinsic rewards enable RL agents to solve complex sparse-reward tasks like Procgen-Maze. Finally, RLeXplore facilitates further research in combined intrinsic rewards, uncovering the potential of such methods.
@@ Appendix A Algorithmic Baselines
ICM [5]. ICM leverages a inverse-forward model to learn the dynamics of the environment and uses the prediction error as the curiosity reward. Specifically, the inverse model inferences the current action $\bm{a}_{t}$ based on the encoded states $\bm{e}_{t}$ and $\bm{e}_{t+1}$, where $\bm{e}=\psi(\bm{s})$ and $\psi(\cdot)$ is an embedding network. Meanwhile, the forward model $f$ predicts the encoded next-state $\bm{e}_{t}$ based on $(\bm{e}_{t},\bm{a}_{t})$. Finally, the intrinsic reward is defined as
@@
RND [4]. RND produces intrinsic rewards via a self-supervised manner, in which a predictor network $\hat{f}$ is trained to approximate a fixed and randomly-initialized target network $\hat{f}$. As a result, the agent is motivated to explore unseen parts of the state space. The intrinsic reward is defined as
@@
Disagreement [9]. Disagreement is variant of ICM that leverages an ensemble of forward models and calculates the intrinsic reward as the variance among these models. Accordingly, the intrinsic reward is defined as
$I_{t}=\mathrm{Var}\{f_{i}(\bm{e}_{t},\bm{a}_{t})\},i=0,...,N$ (5)
NGU [7]. NGU is a mixed intrinsic reward approach that combines global and episodic exploration and the first algorithm to achieve non-zero rewards in the game of Pitfall! without using demonstrations or hand-crafted features. The intrinsic reward is defined as
@@
PseudoCounts [7]. Pseudo-counts has been widely used in count-based exploration approaches [17, 19] with diverse implementations like neural density models. In this paper, we follow NGU [7] that computes pseudo-counts via $k$-nearest neighbor estimation, which is highly efficient and can be applied to arbitrary task. Given the encoded observations $\{\bm{e}_{0},\dots,\bm{e}_{T-1}\}$ visited in the an episode, we have
@@
RIDE [6]. RIDE is designed based on ICM that learns the dynamics of the environment and rewards significant state changes. Accordingly, the intrinsic reward is defined as
@@
RE3 [35]. RE3 is an information theory-based and computation-efficient exploration approach, which aims to maximize the Shannon entropy of the state visiting distribution. In particular, RE3 leverages a random and fixed neural network to encode the state space and employs a $k$-nearest neighbor estimator to estimate the entropy efficiently. Then the estimated entropy is transformed into particle-base intrinsic rewards. Specifically, the intrinsic reward is defined as
@@
E3B [8]. E3B provides a generalization of count-based rewards to continuous spaces using an elliptical bonus. E3B learns a representation mapping from observations to a latent space (e.g., using inverse dynamics). At each episode, the sequence of latent observations parameterize an ellipsoid which is used to measure the novelty of the subsequent observations. In tabular settings, the E3B ellipsoid reduces to the table of inverse state-visitation frequencies [8]. The reward of a new observation is computed as:
$I_{t}=f(\bm{e}_{t})^{T}C_{t-1}f(\bm{e}_{t})$ (11)
where $f$ is the learned representation mapping and $C_{t-1}$ is the episodic ellipsoid.
@@ Appendix B Experimental Settings
B.1 Baselines
We designed the following settings for the baseline experiments, and all the subsequent RQs were progressively adjusted based on the baselines. Moreover, all the experiments are performed using the proximal policy optimization (PPO) [29] implementation from RLLTE [38].
@@ B.1 Baselines
Observation normalization image = image / 255.0
@@
Table 2: PPO hyperparameters for SuperMarioBros and Procgen games.
Hyperparameter SuperMarioBros Procgen
Observation downsampling (84, 84) (64, 64)
Stacked frames No No
Environment steps 10000000 25000000
Episode steps 128 256
Number of workers 1 1
Environments per worker 8 64
Optimizer Adam Adam
Learning rate 2.5e-4 5e-4
GAE coefficient 0.95 0.95
Action entropy coefficient 0.01 0.01
Value loss coefficient 0.5 0.5
Value clip range 0.1 0.2
Max gradient norm 0.5 0.5
Epochs per rollout 4 3
Batch size 256 2048
Discount factor 0.99 0.999
B.2 Details of RQs
Table 3: Details of candidates for all RQs, where $\bm{\mathrm{I}}$ is a batch of intrinsic rewards.
@@ B.2 Details of RQs
RQ 1 Min-Max image = image / 255.0
RMS $\mathrm{image}=\mathrm{Clip}\left(\frac{\mathrm{image}-\mathrm{running\>mean}}{\mathrm{running\>std.}},-5.0,5.0\right)$
RQ 2 Vanilla $\bm{\mathrm{I}}=\bm{\mathrm{I}}$
@@
RQ 3 0.1 Use 10% of the samples to update the intrinsic reward module.
@@
RQ 4 Vanilla Fill the input tensor with values drawn from the uniform distribution.
@@
RQ 5 Vanilla Policy network with only convolutional and linear layers.
@@
Table 3 illustrates the details of the candidates for all RQs. For RQ 1-5, we designed the experiments sequentially and modified the configuration for each intrinsic reward based on the best results of previous RQs. For instance, experiments of RQ 1 only change the technique of observation normalization, and RQ 2 will use the best observation normalization method for each reward module obtained in RQ 1. Likewise, RQ 3 will follow the best results obtained in RQ 1-2 and only change the proportion of samples used for model update. However, we kept using the baselines settings for each reward in RQ 8 to explore the most original performance of the mixed intrinsic rewards.
B.3 Best Configurations
Table 4: The best configurations for each intrinsic reward on SuperMarioBros games.
@@ B.3 Best Configurations
PseudoCounts Min-Max 0.5 Orthogonal
ICM RMS 1.0 Orthogonal
RND Vanilla 0.1 Orthogonal
E3B Min-Max 0.1 Orthogonal
RIDE Min-Max 0.1 Default
RE3 Vanilla N/A Orthogonal
NGU Min-Max 0.1 Orthogonal
Disagreement Min-Max 1.0 Orthogonal
@@ Appendix C Usage Examples
C.1 Workflow of RLeXplore
@@
C.2 RLeXplore with Stable-Baselines3
@@
C.3 RLeXplore with CleanRL
@@
Appendix D Learning Curves

### v2

@@ 6 Conclusion
Our work introduces RLeXplore, a comprehensive open-source repository that not only implements state-of-the-art intrinsic rewards but also provides a systematic evaluation framework for understanding their impact on agent performance. Our results show that with RLeXplore, RL agents can learn emergent behaviours autonomously, solving multiple levels of SuperMarioBros without task rewards. Additionally, we show that intrinsic rewards enable RL agents to obtain great performance on complex sparse-reward tasks like Procgen-Maze, MiniGrid, the ALE-5 hard-exploration tasks and Ant-UMaze. Finally, RLeXplore facilitates further research in mixed intrinsic rewards (Henaff et al., 2023), uncovering the potential of such methods.
@@
RLeXplore is designed to benchmark end-to-end intrinsic reward methods. These end-to-end methods are more commonly used and under-evaluated by the community. Skill-based algorithms, which typically involve separate phases for skill discovery and skill learning, are more complex and left for future work. For an alternative perspective that includes skill-based approaches, we refer readers to the unsupervised RL benchmark by Laskin et al. (2021). Additionally, RLeXplore was designed with accessibility in mind, ensuring that the implemented methods can be run on standard computational resources by any researcher. To maintain this accessibility, we have not included more complex and potentially powerful methods like BYOL-Explore (Guo et al., 2022) or RECODE (Kapturowski et al., ). These methods are not open-source and have been optimized exclusively with non-open-source RL algorithms, which further limits their integration into RLeXplore.
Acknowledgments
This work is funded, in part, by HKSAR RGC under Grant No. PolyU 15224823, the Guangdong Basic and Applied Basic Research Foundation under Grant No. 2024A1515011524, the NSFC under Grant No. 62302246, the ZJNSFC under Grant No. LQ23F010008, and the Ningbo under Grants No. 2023Z237 & 2024Z284 & 2024Z289 & 2023CX050011 & 2025Z038. We thank the high-performance computing center at Eastern Institute of Technology and Ningbo Institute of Digital Twin for providing the computing resources. We also want to acknowledge funding support from NSERC and CIFAR, and compute support from Digital Research Alliance of Canada, Mila IDT and Nvidia.
@@ Appendix A Algorithmic Baselines
ICM (Pathak et al., 2017). ICM leverages an inverse-forward model to learn the dynamics of the environment and uses the prediction error as the curiosity reward. Specifically, the inverse model inferences the current action $\bm{a}_{t}$ based on the encoded states $\bm{e}_{t}$ and $\bm{e}_{t+1}$, where $\bm{e}=\psi(\bm{s})$ and $\psi(\cdot)$ is an embedding network. Meanwhile, the forward model $f$ predicts the encoded next-state $\bm{e}_{t}$ based on $(\bm{e}_{t},\bm{a}_{t})$. Finally, the intrinsic reward is defined as
@@
RND (Burda et al., 2019b). RND produces intrinsic rewards via a self-supervised manner, in which a predictor network $\hat{f}$ is trained to approximate a fixed and randomly-initialized target network $\hat{f}$. As a result, the agent is motivated to explore unseen parts of the state space. The intrinsic reward is defined as
@@
Disagreement (Pathak et al., 2019). Disagreement is a variant of ICM that leverages an ensemble of forward models and calculates the intrinsic reward as the variance among these models. Accordingly, the intrinsic reward is defined as
$I_{t}=\mathrm{Var}\{f_{i}(\bm{e}_{t},\bm{a}_{t})\},i=0,\dots,N.$ (6)
NGU (Badia et al., 2020). NGU is a mixed intrinsic reward approach that combines global and episodic exploration and the first method to achieve non-zero rewards in the game of Pitfall! without using demonstrations or hand-crafted features. The intrinsic reward is defined as
@@
PseudoCounts (Badia et al., 2020). Pseudo-counts has been widely used in count-based exploration approaches (Bellemare et al., 2016; Ostrovski et al., 2017) with diverse implementations like neural density models. In this paper, we follow NGU (Badia et al., 2020) that computes pseudo-counts via $k$-nearest neighbor estimation, which is highly efficient and can be applied to arbitrary tasks. Given the encoded observations $\{\bm{e}_{0},\dots,\bm{e}_{T-1}\}$ visited in the an episode, we have
@@
RIDE (Raileanu & Rocktäschel, 2020). RIDE is designed based on ICM that learns the dynamics of the environment and rewards significant state changes. Accordingly, the intrinsic reward is defined as
@@
RE3 (Seo et al., 2021). RE3 is an information theory-based and computation-efficient exploration approach that aims to maximize the Shannon entropy of the state visiting distribution. In particular, RE3 leverages a random and fixed neural network to encode the state space and employs a $k$-nearest neighbor estimator to estimate the entropy efficiently. Then, the estimated entropy is transformed into particle-based intrinsic rewards. Specifically, the intrinsic reward is defined as
@@
E3B (Henaff et al., 2022). E3B provides a generalization of count-based rewards to continuous spaces. E3B learns a representation mapping from observations to a latent space (e.g., using inverse dynamics). At each episode, the sequence of latent observations parameterizes an ellipsoid (Li et al., 2010; Auer, 2002; Dani et al., 2008), which is used to measure the novelty of the subsequent observations. In tabular settings, the E3B ellipsoid reduces to the table of inverse state-visitation frequencies (Henaff et al., 2022). Given a feature encoding $f$, at each time step $t$ of the episode the elliptical bonus $I_{t}$ is defined as follows:
$I_{t}=f(\bm{s}_{t})^{T}C_{t-1}f(\bm{s}_{t}),$ (12)
$C_{t-1}=\sum_{i=1}^{t-1}f(\bm{s}_{i})f(\bm{s}_{i})^{T}+\lambda\mathbf{I},$ (13)
where $f$ is the learned representation mapping, $C_{t-1}$ is the episodic ellipsoid (Henaff et al., 2022), $\lambda$ is a scalar coefficient, and $\mathbf{I}$ is the identity matrix.
@@ Appendix B Experimental Settings
B.1 Benchmark Selection
We evaluate the RLeXplore framework on multiple recognized benchmarks, which are specifically designed to evaluate the exploration capability of RL agents. We select SuperMarioBros (Kauten, 2018), MiniGrid (Chevalier-Boisvert et al., 2023), Procgen (Cobbe et al., 2020), Arcade learning environment (ALE) (Bellemare et al., 2013), and Gymnasium-Robotics (de Lazcano et al., 2024) for our experiments, which sufficiently spans the existing RL benchmarks. Table 3 provides the details of these selected environments, including their observation, action, and reward spaces.
Table 3: Details of the environments used in our experiments.
Benchmark Environment Observation Space Action Space Reward Space
SuperMarioBros World1-Stage1-v3 Box(0, 255, (240, 256, 3)) Discrete(7) N/A
SuperMarioBros RandomStages-v3 Box(0, 255, (240, 256, 3)) Discrete(7) N/A
MiniGrid DoorKey-16×16-v0 Box(0, 255, (3, 7, 7)) Discrete(7) Sparse
MiniGrid DoorKey-8×8 Box(0, 255, (3, 7, 7)) Discrete(7) Sparse
MiniGrid KeyCorridorS8R5-v0 Box(0, 255, (3, 7, 7)) Discrete(7) Sparse
MiniGrid KeyCorridorS9R6-v0 Box(0, 255, (3, 7, 7)) Discrete(7) Sparse
MiniGrid KeyCorridorS10R7-v0 Box(0, 255, (3, 7, 7)) Discrete(7) Sparse
MiniGrid MultiRoom-N7-S8-v0 Box(0, 255, (3, 7, 7)) Discrete(7) Sparse
MiniGrid MultiRoom-N10-S10-v0 Box(0, 255, (3, 7, 7)) Discrete(7) Sparse
MiniGrid MultiRoom-N12-S10-v0 Box(0, 255, (3, 7, 7)) Discrete(7) Sparse
MiniGrid LockedRoom-v0 Box(0, 255, (3, 7, 7)) Discrete(7) Sparse
MiniGrid Dynamic-Obstacles-16x16-v0 Box(0, 255, (3, 7, 7)) Discrete(3) Sparse
Procgen Maze, num_levels=200 Box(0, 255, (3, 64, 64)) Discrete(15) Sparse
Procgen Maze, num_levels=0 Box(0, 255, (3, 64, 64)) Discrete(15) Sparse
ALE Gravitar Box(0, 255, (4, 84, 84)) Discrete(18) Dense
ALE MontezumaRevenge Box(0, 255, (4, 84, 84)) Discrete(18) Dense
ALE PrivateEye Box(0, 255, (4, 84, 84)) Discrete(18) Dense
ALE Seaquest Box(0, 255, (4, 84, 84)) Discrete(18) Dense
ALE Venture Box(0, 255, (4, 84, 84)) Discrete(18) Dense
Gymnasium-Robotics AntMaze_UMaze-v5 Box(-inf, inf, (27,)) Box(-1, 1, (8,)) Sparse
B.2 Baselines
We designed the following settings for the baseline experiments, and all the subsequent questions were adjusted based on the baselines. Moreover, all the experiments are performed using the proximal policy optimization (PPO) (Schulman et al., 2017) implementation from RLLTE (Yuan et al., 2023).
@@ B.2 Baselines
Observation normalization RMS
@@
B.3 Details of Questions
Table 5 illustrates the details of the candidates for all questions.
Table 5: Details of candidates for all questions, where $\bm{\mathrm{I}}$ is a batch of intrinsic rewards.
@@ B.3 Details of Questions
Q1 Vanilla obs. = obs. / 255.0, only for image-based observations, else obs. = obs..
RMS $\mathrm{obs.}=\mathrm{Clip}\left(\frac{\mathrm{obs.}-\mathrm{running\>mean}}{\mathrm{running\>std.}},-5.0,5.0\right)$
Q2 Vanilla $\bm{\mathrm{I}}=\bm{\mathrm{I}}$
@@
Q3 0.01 Use 1% of the samples to update the intrinsic reward module.
0.1 Use 10% of the samples to update the intrinsic reward module.
@@
Q4 Vanilla Fill the input tensor with values drawn from the uniform distribution.
@@
Q5 Vanilla Policy network with only convolutional and linear layers.
@@
B.4 Best Configurations
Table 6: The best configurations for each intrinsic reward on SuperMarioBros.
@@ B.4 Best Configurations
ICM Min-Max Vanilla 1.0 Default
RND RMS Vanilla 1.0 Orthogonal
Disagreement RMS Vanilla 0.5 Default
NGU RMS Min-Max 0.1 Default
PseudoCounts RMS Min-Max 0.01 Default
RIDE RMS Min-Max 0.01 Default
RE3 Min-Max Min-Max N/A Default
E3B Min-Max Min-Max 0.1 Orthogonal
Table 7: The best configurations for each intrinsic reward on MiniGrid-DoorKey-16×16.
ICM RMS RMS 1.0 Orthogonal
RND RMS Vanilla 0.5 Orthogonal
Disagreement Vanilla Min-Max 0.5 Default
NGU RMS RMS 0.01 Orthogonal
PseudoCounts RMS Min-Max 1.0 Orthogonal
RIDE RMS Min-Max 1.0 Orthogonal
RE3 RMS Min-Max N/A Orthogonal
E3B RMS RMS 1.0 Orthogonal
Table 8: PPO hyperparameters for SuperMarioBros, MiniGrid, and Procgen games. These remain fixed for all experiments.
Hyperparameter SuperMarioBros MiniGrid Procgen
Observation downsampling (84, 84) (7, 7) (64, 64)
Observation normalization / 255. No / 255.
Reward normalization No No No
Weight initialization Orthogonal Orthogonal Orthogonal
LSTM No No No
Stacked frames No No No
Environment steps 10000000 10000000 25000000
Episode steps 128 32 256
Number of workers 1 1 1
Environments per worker 8 256 64
Optimizer Adam Adam Adam
Learning rate 2.5e-4 2.5e-4 5e-4
GAE coefficient 0.95 0.95 0.95
Action entropy coefficient 0.01 0.01 0.01
Value loss coefficient 0.5 0.5 0.5
Value clip range 0.1 0.1 0.2
Max gradient norm 0.5 0.5 0.5
Epochs per rollout 4 4 3
Batch size 256 1024 2048
Discount factor 0.99 0.99 0.999
@@ Appendix C Usage Examples
C.1 API Compatiblity
The following table provides a detailed algorithm and environment compatibility of the implemented intrinsic rewards. Since NGU, PseudoCounts, RIDE, and E3B require an episode memory for the reward computation, when combined with off-policy RL algorithms, they can only work in a non-vectorized environment currently. Therefore, these reward modules are marked with a * symbol.
Table 9: Algorithm and environment compatibility of the RLeXplore framework.
Algorithm
Image-based
Observation
State-based
Observation
Continuous
Action
Discrete
Action
On-policy
RL
Off-policy
RL
ICM
RND
Disagreement
NGU
PseudoCounts
RIDE
RE3
E3B
C.2 Workflow of RLeXplore
@@
C.3 Mixed Intrinsic Reward
@@
C.4 RLeXplore with Stable-Baselines3
@@
C.5 RLeXplore with CleanRL
[…]
C.6 Implementing New Intrinsic Reward Modules
In RLeXplore, all intrinsic reward methods inherit from a base reward class that requires two functions to be implemented: compute() and update(). The compute() function processes a batch of on-policy trajectories to calculate intrinsic rewards and is automatically called prior to the PPO update, while the update() function uses the same trajectories to update the associated modules. To integrate a new intrinsic reward method, users only need to create a new script that inherits from the base reward class and implements these two functions. Moreover, many pre-defined network modules (e.g., Atari CNN, ResNet CNN) are readily available for import, allowing users to use the currently implemented intrinsic rewards as templates for their own implementations.
Appendix D Comparative Analysis of Intrinsic Reward Implementations
This section provides a detailed comparative analysis of our intrinsic reward implementations in the RLeXplore framework against other publicly available implementations. The results are compiled in tables for different environments to demonstrate the performance of each algorithm. We cite the works from which we obtained the original results in each of the tables, and we provide our results by averaging the performance of the last 100 training episodes over 3 seeds.
D.1 SuperMarioBros (Only Intrinsic Rewards)
Table 10: Comparison of % of level completed in SuperMarioBros without task rewards.
Algorithm % of Level Completed (10M Steps) % of Level Completed (1M Steps)
(Original) RIDE - 23%
(Original) ICM 30% -
(RLeXplore) RIDE 100% 50%
(RLeXplore) ICM 30% 2%
The percentage of the level completed is computed by dividing the episode return by 3,000, which corresponds to the maximum reward that can be obtained in SuperMarioBros-1-1 (if the agent solves the level without wasting time). Note that in Figure 3, we divide this quantity by 100 and show a maximum reward of 30.
Note that our implementation of ICM reproduces the results reported in the original paper in Mario (Pathak et al., 2017), and our implementation of RIDE further outperforms the original implementation.
D.2 MiniGrid-DoorKey-16×16 (Extrinsic + Intrinsic Rewards)
Table 11: Episode returns in MiniGrid-DoorKey-16×16 with extrinsic and intrinsic rewards.
Algorithm Episode Return (10M Steps)
(Original) RIDE (Zhang et al., 2020) 0.25
(Original) ICM (Zhang et al., 2020) 0.0
(Original) RND (Zhang et al., 2020) 0.0
(Original) IMPALA (Zhang et al., 2020) 0.0
(RLeXplore) PPO 0.37
(RLeXplore) ICM 0.6
(RLeXplore) RND 0.6
(RLeXplore) RIDE 0.12
Using the implementations in RLeXplore we obtain significantly better performance in the same tasks and with the same algorithms.
D.3 MiniGrid-DoorKey-8×8 (Extrinsic + Intrinsic Rewards)
We also evaluate our implementations in MiniGrid-DoorKey-8×8 with a budget of 1M environment steps to be able to compare to the original results reported in (Seo et al., 2021).
Table 12: Episode returns in MiniGrid-DoorKey-8×8 with 1M environment steps.
Algorithm Episode Return (1M Steps)
(Original) RE3 (Seo et al., 2021) 0.5
(Original) RND (Seo et al., 2021) 0.0
(Original) ICM (Seo et al., 2021) 0.2
(Original) A2C (Seo et al., 2021) 0.0
(RLeXplore) RE3 0.95
(RLeXplore) RND 0.82
(RLeXplore) ICM 0.83
(RLeXplore) PPO 0.22
Importantly, we reproduce the results reported in (Seo et al., 2021) very accurately, showing that RE3 can provide more sample-efficient exploration in this domain, compared to RND and ICM. Still, our implementations of RE3 and ICM achieve even better performance than the original ones.
Figure 9: Using RLeXplore in MiniGrid-DoorKey-8×8, we are able to not only reproduce the conclusions obtained in previous work (Seo et al., 2021) regarding the capabilities of RE3 compared to ICM and RND, but we also generally achieve better performance, hence providing stronger baselines to the RL community.
D.4 Procgen - 200 Mazes (Extrinsic + Intrinsic Rewards)
Table 13: Performance comparison in Procgen - 200 Mazes with 25M training steps.
Algorithm Procgen - 200 Mazes (25M Steps)
(Original) E3B (Castanyer et al., 2023) 3.0
(Original) ICM (Castanyer et al., 2023) 2.5
(Original) RND (Castanyer et al., 2023) 1.7
(RLeXplore) E3B 4.1
(RLeXplore) ICM 5.9
(RLeXplore) RND 5.0
D.5 ALE-5 (Extrinsic + Intrinsic Rewards)
In this section, we present the evaluation results of the intrinsic reward methods on a set of ALE games known for their challenging exploration requirements. These "hard-exploration" games, including Gravitar, Montezuma’s Revenge, Private Eye, Seaquest, and Venture, serve as a benchmark for testing the effectiveness of intrinsic rewards in aiding exploration and improving agent performance.
We observe that while intrinsic rewards lead to a decline in performance in Gravitar, they generally provide substantial benefits, particularly in environments where exploration is difficult. For example, in Seaquest, the use of intrinsic rewards enables algorithms to significantly outperform the extrinsic agent, which ranks among the lowest.
Note that we do not compare these results to other works because evaluation settings differ significantly between papers. For instance, in our case, we used sticky actions with a probability of 0.25%, which makes the exploration problem more difficult, and it is not always used. Also, we trained our agents for 25M steps instead of the standard 200M due to computational constraints. Still, our results provide evidence that intrinsic rewards are generally helpful in achieving better episode returns in hard-exploration environments.
Table 14: Mean performance across different environments for each algorithm, averaged over 3 seeds after 25M environment steps. Results are averaged over the last 100 episodes of training. In Gravitar, intrinsic rewards appear to hinder the performance of the extrinsic agent, whereas, in other environments, they significantly enhance performance. Notably, in Seaquest, the extrinsic agent ranks among the lowest, highlighting the benefit of intrinsic rewards. All experiments were conducted using sticky actions with a repeat probability of 0.25.
Algorithm Gravitar MontezumaRevenge PrivateEye Seaquest Venture
Extrinsic 1060.19 42.83 88.37 942.37 391.73
Disagreement 689.12 0.00 33.23 6577.03 468.43
E3B 503.43 0.50 66.23 8690.65 0.80
ICM 194.71 31.14 -27.50 2626.13 0.54
PseudoCounts 295.49 0.00 1076.74 668.96 1.03
RE3 130.00 2.68 312.72 864.60 0.06
RIDE 452.53 0.00 -1.40 1024.39 404.81
RND 835.57 160.22 45.85 5989.06 544.73
Table 15: Aggregated performance comparison of mean episode return between RLeXplore and original implementations on the ALE-5 benchmark. The results show average returns across five Atari games. Since the original implementations of R2D2 and NGU are not publicly available, we report their published results obtained with 35B environment steps, while our experiments were conducted with 100 million frames.
Algorithm RL Algorithm Frames Mean Episode Return
Extrinsic - PPO PPO 100M 0.5k
Extrinsic - R2D2 (Kapturowski et al., 2018) R2D2 35B 210k
(RLeXplore) Disagreement PPO 100M 1.5k
(RLeXplore) NGU PPO 100M 0.8k
(RLeXplore) PseudoCounts PPO 100M 0.4k
(Original) Disagreement (Pathak et al., 2019) PPO 100M 0.1k
(Original) NGU (Badia et al., 2020) R2D2 35B 225k
Table 15 further provides a direct performance comparison between our RLeXplore implementations and the original ones reported in the literature. Since no public codebase or dataset is available for NGU, we extracted its baseline performance numbers directly from the paper. Despite operating under a more limited training budget, our NGU implementation still achieves competitive performance compared to the published results. For Disagreement, we used its official repository and trained the model for 1M frames. The scores we obtained match the results reported in (Pathak et al., 2019), confirming that our implementation faithfully reproduces the expected performance. Overall, the improved performance of RLeXplore’s implementations is largely attributable to the extra effort put into fine-tuning low-level details, which helps to optimize the interplay of the various components.
D.6 Comparison with Other Projects
Table 16: Details on official implementations of the included intrinsic rewards. Decoupled: Did the code decouple the intrinsic reward modules from the RL optimization algorithms, which can be directly reused in other projects?
Reward
Official
Repository
ML
framework
Backbone
RL algorithm
Supported
Tasks
Decoupled
ICM Repository Tensorflow A3C
SuperMarioBros,
VizDoom
RND Repository Tensorflow PPO ALE
Disagreement Repository Tensorflow PPO
SuperMarioBros,
ALE, Maze
NGU N/A N/A N/A N/A N/A
PseudoCounts from NGU N/A N/A N/A N/A
RIDE Repository PyTorch IMPALA MiniGrid
RE3 Repository PyTorch A2C, Dreamer, RAD
DMControl,
MiniGrid
E3B Repository PyTorch IMPALA
MiniHack,
VizDoom
Table 16 illustrates the details of official implementations of the included intrinsic rewards in RLeXplore. It is natural to find that they are implemented (1) in different codebases with (2) different libraries (e.g., PyTorch vs Tensorflow), (3) using different RL algorithms (PPO, IMPALA, A3C, A2C), and (4) supporting different environments (ALE, Mario, MiniGrid, DMC). These details further motivate the development of a unified framework for training RL agents with intrinsic rewards under standardized conditions and reinforce our motivation to develop RLeXplore.
Furthermore, we provide a comparison of the advantages of other popular codebases for training RL agents with intrinsic rewards in terms of the number of intrinsic reward methods implemented, their modularity and ability to reuse components between RL libraries easily, their documentation, and the number of experiments provided. As compared to other existing projects, RLeXplore offers a distinctive advantage by providing a more unified and standardized approach to training RL agents with intrinsic rewards. It allows users to easily swap intrinsic reward modules regardless of RL libraries, which promotes reproducibility and consistency across different research works. Finally, RLeXplore is evaluated on a wide range of benchmarks with over 1,000 experiments, ensuring its reliability and robustness across various scenarios.
Table 17: Comparison between RLeXplore and other reported libraries of intrinsic rewards. Note that we focus on the intrinsic reward methods that are implemented. For instance, CleanRL has many implementations of different RL algorithms, but RND is the only supported intrinsic reward.
Framework
ML
Framework
Number of
Algorithms
Plug & Play Documentation
Benchmark
Results
CleanRL PyTorch 1
1 task,
1 experiments
DI-Engine PyTorch 3
5 tasks,
19 experiments
rllib TensorFlow 2 N/A
RLeXplore PyTorch 8
17 tasks,
over 2000 experiments
Appendix E Learning Curves
E.1 Q1